2 6 (1 point) Find the orthogonal projection of v 14 onto the subspace V of Rspanned by 6 and 8 projv(v) =
Answer #1

ę w E V2= 14 -8 Orthogonal projection of & onto V is » vov V, + Vio 2. Vz ve Proj, (b) = 1+36+4 = 41 -2] HiV= [- -6 -2 -6 V.W